# Understanding Chain-of-Thought in LLMs through Information Theory

### v1

@@ 3.3 Testing for unidentifiability using information gain
Let $\mathcal{I}(X;Y\mid Z)$ denote the mutual information between random variables $X$ and $Y$ conditional on $Z$. Then,
@@
To estimate the information gain in (3) using Proposition 3.4, we train a separate LLM, which we refer to as the supervisor model $g_{\textup{sup}}$. This model takes as input the model’s CoT reasoning up to any given intermediate step $t$, $X^{M}_{t}$, and is fine-tuned to directly predict the ground truth final output $Y$. In this way $g_{\textup{sup}}(X^{M}_{t})$ approximates the conditional distribution $p(Y\mid X^{M}_{t})$. Then, the quantity $\mathbb{E}[\log p(Y\mid X^{M}_{j})]$ can be estimated using the negative cross-entropy loss for predicting $Y$, i.e.
$\mathbb{E}[\log p(Y\mid X^{M}_{j})]\approx\mathbb{E}[\log\hat{p}(Y\mid X^{M}_{j})]=-\mathbb{E}[l_{\textup{CE}}\left(Y,g_{\textup{sup}}(X^{M}_{j})\right)],$
@@
$\underbrace{\mathbb{E}[\log p(Y\mid X^{M}_{j})]-\mathbb{E}[\log p(Y\mid X^{M}_{j-1})]}_{\text{Information gain}}\approx\mathbb{E}[l_{\textup{CE}}(Y,g_{\textup{sup}}(X^{M}_{j-1}))]-\mathbb{E}[l_{\textup{CE}}(Y,g_{\textup{sup}}(X^{M}_{j}))].$ (4)
Summary: The information gain (IG) between steps $j$ and $j-1$ reflects how much relevant information step $j$ contributes towards predicting $Y$. If task $\lambda_{j}$ is executed correctly, this gain is positive, as indicated by a decrease in the cross-entropy loss. Conversely, if step $j$ does not provide additional information, the loss remains unchanged. This can be interpreted as the conditional mutual information between $X^{M}_{j}$ and $Y$, conditioned on $X^{M}_{j-1}$. Positive information gain suggests step $j$ adds new insight about $Y$, while no gain indicates no added information. Training details for the supervisor model are in Appendix B.1.3.
Remark on sample-wise information gain     While conditional mutual information provides an aggregate measure of information gain for a sub-task in a dataset, it may also be desirable to obtain an analogous measure of sub-task correctness for individual CoT instances. This could be useful, for example, in detecting which step went wrong for a given prompt. Our notion of information gain can be extended to this sample-wise setting by instead considering the following difference
@@
Intuitively, if step $j$ in the model’s CoT is correct, the model should become more confident in the ground truth output $Y$ being the correct final answer. Therefore, the difference above should be positive. Alternatively, if step $j$ is wrong, the model’s confidence regarding the true output $Y$ should not increase and the above difference should not be positive. From now on, we refer to the difference in (5) as sample-wise information gain at step $j$.
@@ 4 Related works
Evaluation of CoT reasoning
Several recent works propose methodologies for evaluating CoT reasoning (Wei et al., 2024; Havrilla et al., 2024; Li et al., 2023; Joshi et al., 2023; Nguyen et al., 2024; Wang et al., 2024a; Yu et al., 2024; Xie et al., 2024). For example, Li et al. (2023) verifies individual steps in a model’s CoT reasoning by generating multiple LLM responses per prompt and comparing correct responses with incorrect ones. Similarly, Wang et al. (2024b, c) use a fine-tuned LLM to decode multiple reasoning paths from each step and check the correctness of these reasoning paths. However, as we show in our experiments, approaches which simply rely on the correctness of the final output are not sound in general and can lead to false positives. Moreover, these solutions may not be plausible for problems of high difficulty where correct LLM responses might be scarce.
Formalising CoT framework
The formalisation of LLM reasoning remains an active area of research. Most notably González and Nori (2023) introduces a formal framework for LLMs and is a key source of inspiration behind our formalism. Additionally, Feng et al. (2023) theoretically examines the expressivity of LLMs with CoT in solving mathematical and decision-making problems, focusing on the transformer architecture’s implications on accuracy. Besides this, Xu et al. (2024) provides a formal definition of hallucinations, but does not consider CoT reasoning specifically.
Reward modelling
One notable line of work known as outcome-based reward models (ORM) (Cobbe et al., 2021; Havrilla et al., 2024; Lightman et al., 2023) predicts the probability of reaching the correct final answer given a model’s intermediate CoT steps. While ORMs do not require demonstrations of correct intermediate steps, we show in Section 5 that this approach is not sound for detecting errors in a model’s CoT reasoning. Another related method is step-wise ORM (SORM) Havrilla et al. (2024) which estimates the probability of an ‘optimal’ model reaching a correct answer, given the CoT reasoning of our model of interest. However, unlike our approach, SORM requires training a model which is larger and more capable than our base model.
Process-based reward modelling (PRMs) (Lightman et al., 2023; Uesato et al., 2022) is an alternative approach which directly predicts the correctness of intermediate CoT reasoning steps. Likewise, various other approaches rely on annotated CoT datasets for benchmarking (Jacovi et al., 2024; Yu et al., 2024; Amini et al., 2019; Liu et al., 2020; Xi et al., 2024; Nguyen et al., 2024; Xie et al., 2024; McLeish et al., 2024). While these benchmarks and methodologies can be valuable for improving LLM reasoning, collecting annotated data can be very costly and is not readily scalable to other tasks. Unlike these methods, our approach computes the information gain at each step, providing a richer measure of LLM performance without requiring any human-annotated CoT data.
@@ 5 Experiments
In this section, we empirically demonstrate the practical utility of our framework. In addition to our proposed method dubbed information gain (denoted by IG), we consider two common baselines that can be used to detect the errors in a model’s CoT reasoning and assume access to only the model’s CoT generations $X_{0},X^{M}_{1},\ldots,X^{M}_{T}$ as well as the correct final answers denoted as $Y$.
Outcome Reward Model (ORM) (Havrilla et al., 2024)
@@
Here, if we observe that this probability of correctness drops significantly after step $t$, i.e. if $f_{\textup{ORM}}(X^{M}_{t})\gg f_{\textup{ORM}}(X^{M}_{t+1})$, this indicates that the model does not apply the task $\lambda_{t+1}$ correctly.
Math-Shepherd (Wang et al., 2024b)
This method quantifies the potential for a given reasoning process $X^{M}_{t}$ by using a ‘completer’ model to generate $N$ completions of each reasoning process starting from step $t$, $\{(X^{M}_{t},X^{M}_{t+1,j},\ldots,X^{M}_{T,j},Y^{M}_{j})\}_{j\leq N}$, where $Y^{M}_{j}$ denotes the final answer reached in the $j$’th completion. Then, we estimate the potential of this step based on the proportion of correct answers among the $N$ completions, denoted by $f_{\textup{MS}}(X^{M}_{t})$ as:
$f_{\textup{MS}}(X^{M}_{t})\coloneqq\sum_{j=1}^{N}\frac{{\mathbbm{1}}(Y^{M}_{j}=Y)}{N}.$ (7)
For a fair comparison we do not assume access to a ‘verifier’ model more capable than our base model and therefore, we use the base model as the completer model in our experiments.
@@ 5.1 Toy data experiments
First, we consider a toy setting where we have full control over the model behaviour on different tasks. Our prompts comprise of an integer vector $Z_{0}\in\mathbb{Z}^{5}$ sampled randomly from a given distribution. The task $\lambda$ comprises 5-steps $\lambda=\lambda_{1}\circ\ldots\circ\lambda_{5}$, where each sub-task $\lambda_{i}$ denotes an operation which transforms a given integer vector $Z_{i-1}\in\mathbb{Z}^{5}$ into another $Z_{i}\in\mathbb{Z}^{5}$. Finally, in this setup, the correct final answer $Y$ is the value of $Z_{5}$. Additional details on the data generating mechanism as well as the sub-tasks are provided in Appendix B.1.
Generating the dataset
To investigate partial unidentifiability for a given task $\lambda_{i}$ we modify the obtained dataset by introducing ‘noise’ at step $i$. In other words, the task $\lambda_{i}$ is applied incorrectly on a subset of the data, whereas all other tasks are always applied correctly. This represents a model which sometimes fails at step $i$ and we use ‘LLMi’ to denote this model in this experiment. We repeat this procedure for all tasks $\lambda_{i}$ for $i\in\{1,\ldots,5\}$ which yields 5 LLMs $\{\textup{LLM}_{1},\ldots,\textup{LLM}_{5}\}$.
To also investigate the robustness of the methods, we introduce a special case in $\textup{LLM}_{3}$. Here, task $\lambda_{3}$ is applied incorrectly if and only if the output after task 2 (i.e., after $\lambda_{2}$) lies in some set $\mathcal{S}$. This choice has been made deliberately to highlight a pitfall of the existing baselines (as we will explain below) and is in contrast to the rest of LLMs where any errors occur at random. In other words, the correctness of task $\lambda_{3}$ is dependent on the output of $\lambda_{2}$. For more details, see Appendix B.1.2.
@@
Figure 3: Heatmaps quantifying the correctness of different sub-tasks for the 5 LLMs under consideration obtained using the different baselines. Here, the red color indicates a significant drop in the plotted metrics and can be seen as an indication of an incorrectly executed sub-task.
5.1.1 Results
Figure 3 shows how the different baselines quantify the correctness of the different tasks for the 5 different LLMs under consideration. This figure only considers samples where the final answer of the LLM was incorrect, i.e. $Y^{M}\neq Y$. For our method (IG), Figure 3(a) shows the information gain across the different steps for each LLM. Likewise, Figure 3(b) presents the results for ORM and shows how the average probability of correctness in (6) changes across the different steps, whereas, for Math-Shepherd, Figure 3(c) shows the proportion of correct completions starting after each step (7). Here, any significant drop in the plotted values indicate an incorrect application of a task.
Information gain accurately quantifies step-wise correctness     We observe that for each LLM the information gain remains positive until we encounter an incorrect reasoning step, at which point it drops to negative values. Therefore, our method can identify the incorrectly executed task for each LLM under consideration. We used a GPT-2 supervisor model to estimate information gain.
Pitfall of the baselines     While the ORM and Math-Shepherd manage to correctly identify the incorrect reasoning steps in most cases, these methods fail to correctly detect erroneous steps for $\textup{LLM}_{3}$. This happens because, in our setup, $\lambda_{3}$ is incorrectly applied if and only if the output after task $\lambda_{2}$ lies in some set $\mathcal{S}$. Therefore, the classifier model can confidently predict the correctness of the final model output at $\lambda_{2}$ by simply checking if the output lies in $\mathcal{S}$. Here, the classifier becomes confident that the final output will be wrong right after $\lambda_{2}$, even though the error occurs at $\lambda_{3}$.
Table 1: Metrics for sample-wise classification of sub-task correctness for $\textup{LLM}_{3}$ using the different baselines.
Method Accuracy $\uparrow$ TPR $\uparrow$ FPR $\downarrow$
@@ 5.1.1 Results
Math-Shepherd 0.60 1.0 1.0
Similarly, when using Math-Shepherd for $\textup{LLM}_{3}$ (with the same model being used as a completer), a completion yields an incorrect final answer if the output after $\lambda_{2}$ lies in $\mathcal{S}$. If this is the case, all completions yield an incorrect final output regardless of which step we begin the completions from. This makes it impossible to accurately identify the step at which $\textup{LLM}_{3}$ goes wrong.
Sample-wise detection     We can also use the different baselines for sample-wise detection of erroneous steps as outlined in Section 3.3. In this setting, for each prompt, we can classify a step as incorrect if a baseline’s metric falls below a threshold. Table 1 shows the results for sample-wise classification of sub-task correctness for $\textup{LLM}_{3}$ using the different baselines (where we chose the best thresholds for each baseline using a held-out dataset). It can be seen that our method yields a significantly higher accuracy and a lower rate of false-positives than the baselines and therefore, is also considerably more reliable for sample-wise detection of errors.
@@ 5.2 Arithmetic operations on LLama-3-8B
Following our toy experiments, we now evaluate our framework in a more realistic setting using the Llama-3-8B model (Dubey et al., 2024). We focus on a simple arithmetic task that involves both multiplication and addition tasks. The goal is to assess the model’s performance on individual operations as well as their combination.
Experimental setup     We sample two integers $x$ and $y$ uniformly from the range $[1,100000)$. The prompt given to the model is structured as follows:
@@
Model accuracy     We observe that the model’s accuracy varies across the three steps:
Step 1 accuracy: 80%, Step 2 accuracy: 98%, Step 3 accuracy: 42%.
Figure 4: The distribution of $(x,y)$ for incorrect samples shows a clear trend: Llama-3-8B struggles to add large and small numbers together (top-left and bottom-right).
Notably, the majority of failures occur in the third step, which involves addition of the previously computed values. We analyzed the distribution of $(x,y)$ values where the model obtains the correct final output. Interestingly, as Figure 4 illustrates, we observed that most errors occur when exactly one of the variables $(x,y)$ is large and the other is small. This suggests that the model’s correctness is highly dependent on the $(x,y)$ values in the prompt, resulting in baselines struggling to identify the erroneous step in the model’s CoT reasoning (as we show below).

### v2

@@ 3.3 Testing for unidentifiability using information-gain
Let $\mathcal{I}(X;Y\mid Z)$ denote the mutual information between $X$ and $Y$ conditional on $Z$. Then,
@@
To estimate the information-gain in (3) using Proposition 3.4, we train a separate LLM, which we refer to as the supervisor model $g_{\textup{sup}}$. This model takes as input the model’s CoT reasoning up to any given intermediate step $t$, $X^{M}_{t}$, and is fine-tuned to directly predict the ground truth final output $Y$. In this way $g_{\textup{sup}}(X^{M}_{t})$ approximates the conditional distribution $p(Y\mid X^{M}_{t})$. Then, the quantity $\mathbb{E}[\log p(Y\mid X^{M}_{j})]$ can be estimated using the negative cross-entropy loss for predicting $Y$, i.e., $\mathbb{E}[\log p(Y\mid X^{M}_{j})]$ is approximately
$\mathbb{E}[\log\hat{p}(Y\mid X^{M}_{j})]=-\mathbb{E}[l_{\textup{CE}}\left(Y,g_{\textup{sup}}(X^{M}_{j})\right)],$
@@
$\underbrace{\mathbb{E}[\log p(Y\mid X^{M}_{j})]-\mathbb{E}[\log p(Y\mid X^{M}_{j-1})]}_{\text{Information-gain}}\approx\mathbb{E}[l_{\textup{CE}}(Y,g_{\textup{sup}}(X^{M}_{j-1}))]-\mathbb{E}[l_{\textup{CE}}(Y,g_{\textup{sup}}(X^{M}_{j}))].$ (4)
Summary    The information-gain (IG) between steps $j$ and $j-1$ reflects how much relevant information step $j$ contributes towards predicting $Y$. If task $\lambda_{j}$ is executed correctly, this gain is positive, as indicated by a decrease in the cross-entropy loss. Conversely, if step $j$ does not provide additional information, the loss remains unchanged. This can be interpreted as the conditional mutual information between $X^{M}_{j}$ and $Y$, conditioned on $X^{M}_{j-1}$. Positive information-gain suggests step $j$ adds new insight about $Y$, while no gain indicates no added information. Training details for the supervisor model are in Appendix 9.1.3.
Remark on sample-wise information-gain     While conditional mutual information provides an aggregate measure of information-gain for a sub-task in a dataset, it may also be desirable to obtain an analogous measure of sub-task correctness for individual CoT instances. This could be useful, for example, in detecting which step is wrong for a given prompt. Our notion of information-gain can be extended to this sample-wise setting, similar to [8], by instead considering the following difference
@@
Intuitively, if step $j$ in the model’s CoT is correct, the model should become more confident in the ground truth output $Y$ being the correct final answer. Therefore, the difference above should be positive. Conversely, if step $j$ is wrong, the model’s confidence regarding $Y$ should not increase, and this difference should be $\leq 0$. From now on, we refer to the difference in (5) as sample-wise information-gain at step $j$.
Remark on O1/R1 style reasoning     Although we present our framework using linear chains-of-thought for clarity, the information-gain metric naturally accommodates the more complex reasoning patterns found in O1/R1-style models. These reasoning models often explore multiple solution paths, backtrack when encountering errors, and dynamically self-correct their approach. In such exploratory settings, our framework remains effective: steps along incorrect reasoning trajectories will exhibit low or negative information gain, indicating they do not contribute meaningfully toward the correct final answer. When the model identifies a more promising path and self-corrects, subsequent steps will show positive information gain, signaling productive progress.
This adaptability to varied reasoning structures is empirically demonstrated in our experiments (Section 5), where we analyse reasoning traces from the MATH/PRM dataset. Steps that human annotators labelled as uninformative or irrelevant consistently show low information gain under our metric, while correct and meaningful steps exhibit high information gain. Thus, our method provides reliable step-wise evaluation regardless of whether the reasoning follows a linear chain or involves branched/exploratory patterns.
We further formalize this remark using our framework in Appendix 8.4.
@@ 4 Related Works
Evaluation of CoT reasoning   Several recent works propose methodologies for evaluating CoT reasoning [32, 11, 16, 15, 21, 28, 37, 35]. For example, [16] verifies individual steps in a model’s CoT reasoning by generating multiple LLM responses per prompt and comparing correct responses with incorrect ones.
Similarly, [29, 30] use a fine-tuned LLM to decode multiple reasoning paths from each step and check the correctness of these reasoning paths. However, as we show in our experiments, approaches which simply rely on the correctness of the final output are not sound in general and can lead to false positives. Moreover, these solutions may not be plausible for problems of high difficulty where correct LLM responses might be scarce.
Formalising CoT framework    The formalisation of LLM reasoning remains an active area of research. Most notably [10] introduces a formal framework for LLMs and is a key source of inspiration behind our formalism. Additionally, [9] theoretically examines the expressivity of LLMs with CoT in solving mathematical and decision-making problems, focusing on the transformer architecture’s implications on accuracy. Besides this, [36] provides a formal definition of hallucinations, but does not consider CoT reasoning specifically.
Reward modelling    Outcome-based reward models (ORM) [6, 11, 18] predict the probability of reaching the correct final answer based on a model’s intermediate CoT steps. While they avoid requiring correct intermediate demonstrations, we show in Section 5 that they are unsound for detecting CoT reasoning errors. Step-wise ORM (SORM) [11] extends ORM by estimating the probability of an ‘optimal’ model reaching a correct answer but requires training a larger, more capable model than the base model.
Process-based reward modelling (PRMs) [18, 27] is an alternative approach which directly predicts the correctness of intermediate CoT reasoning steps. Likewise, various other approaches rely on annotated CoT datasets for benchmarking [14, 37, 1, 19, 34, 21, 35, 20]. While these benchmarks and methodologies aid LLM reasoning, collecting annotated data is costly and not easily scalable. In contrast, our approach evaluates an LLM’s CoT reasoning without human-annotated CoT data.
@@ 5 Experiments
In this section, we demonstrate our framework’s utility, dubbed Information-Gain (IG) and compare against two baselines for detecting errors in a model’s CoT reasoning. Here we assume access only to the model’s CoT generations $X_{0},X^{M}_{1},\ldots,X^{M}_{T}$ and correct final answers $Y$.
Outcome Reward Model (ORM) [11]
@@
Here, if we observe that this probability of correctness drops significantly after step $t$, i.e. $f_{\textup{ORM}}(X^{M}_{t})\gg f_{\textup{ORM}}(X^{M}_{t+1})$, this indicates that the model applies task $\lambda_{t+1}$ incorrectly.
Math-Shepherd (MS) [29]
This method quantifies the potential for a given reasoning process $X^{M}_{t}$ by using a ‘completer’ model to generate $N$ completions of each reasoning process starting from step $t$, $\{(X^{M}_{t},X^{M}_{t+1,j},\ldots,X^{M}_{T,j},Y^{M}_{j})\}_{j\leq N}$, where $Y^{M}_{j}$ denotes the final answer reached in the $j$’th completion. Then, we estimate the potential of this step based on the proportion of correct answers among the $N$ completions:
$f_{\textup{MS}}(X^{M}_{t})\coloneqq\sum_{j=1}^{N}{\mathbbm{1}}(Y^{M}_{j}=Y)/N.$ (7)
For a fair comparison, we do not assume access to a ‘verifier’ model more capable than our base model. Therefore, we use the base model as the completer model in our experiments.
@@ 5.1 Toy data experiments
We first consider a toy setting where we control model behaviour across tasks. Prompts consist of an integer vector $Z_{0}\in\mathbb{Z}^{5}$ sampled from a given distribution. The task $\lambda$ comprises five steps, $\lambda=\lambda_{1}\circ\ldots\circ\lambda_{5}$, where each sub-task $\lambda_{i}$ transforms $Z_{i-1}\in\mathbb{Z}^{5}$ into $Z_{i}\in\mathbb{Z}^{5}$. The correct final answer $Y$ is $Z_{5}$. Additional details on data generation and sub-tasks are in Appendix 9.1.
Generating the dataset    To investigate partial unidentifiability for a given task $\lambda_{i}$ we modify the obtained dataset by introducing ‘noise’ at step $i$. In other words, the task $\lambda_{i}$ is applied incorrectly on a subset of the data, whereas all other tasks are always applied correctly. This represents a model which sometimes fails at step $i$ and we use ‘LLMi’ to denote this model in this experiment. We repeat this procedure for all tasks $\lambda_{i}$ for $i\in\{1,\ldots,5\}$ which yields 5 LLMs $\{\textup{LLM}_{1},\ldots,\textup{LLM}_{5}\}$.
To assess robustness, we introduce a special case in $\textup{LLM}_{3}$, where task $\lambda_{3}$ is applied incorrectly iff the output after $\lambda_{2}$ falls in a set $\mathcal{S}$. This deliberate choice highlights a pitfall of existing baselines and contrasts with other LLMs, where errors occur randomly. In other words, $\lambda_{3}$’s correctness depends on $\lambda_{2}$’s output. For details, see Appendix 9.1.2.
5.1.1 Results
Figure 3 shows how the different baselines quantify the correctness of the different tasks for the 5 different LLMs under consideration. This figure only considers samples where the final answer of the LLM was incorrect, i.e. $Y^{M}\neq Y$. For our method (IG), Figure 3(a) shows the information-gain across the different steps for each LLM. Likewise, Figure 3(b) presents the results for ORM and shows how the average probability of correctness in (6) changes across the different steps, whereas, for Math-Shepherd, Figure 3(c) shows the proportion of correct completions starting after each step (7). Here, any significant drop in the plotted values indicate an incorrect application of a task.
Information-gain rightly quantifies step-wise correctness     We observe that the information-gain remains positive for each LLM until we encounter an incorrect reasoning step, at which point it drops to negative values. Therefore, our method can identify the incorrectly executed task for each LLM under consideration. We used a GPT-2 supervisor model to estimate information-gain.
Pitfall of the baselines     While ORM and Math-Shepherd usually identify incorrect reasoning steps, they fail for $\textup{LLM}_{3}$. This is because $\lambda_{3}$ is misapplied iff the output after $\lambda_{2}$ lies in $\mathcal{S}$. Thus, the classifier can predict the final output’s correctness at $\lambda_{2}$ by checking if $Z_{2}$ lies in $\mathcal{S}$, leading to overconfidence in error detection at $\lambda_{2}$ instead of $\lambda_{3}$.
@@ 5.1.1 Results
Figure 3: Heatmaps (left) quantifying the correctness of different sub-tasks for the 5 LLMs using the different baselines, and the associated classification metrics (right). Red color in the heatmaps indicates a significant drop in the plotted metrics (an incorrectly executed sub-task).
Similarly, with Math-Shepherd for $\textup{LLM}_{3}$ (using the same model as a completer), a completion is incorrect if the output after $\lambda_{2}$ lies in $\mathcal{S}$. Here, all completions fail, regardless of the starting step, making it impossible to pinpoint where $\textup{LLM}_{3}$ goes wrong.
Table 1: Sample-wise classification of a sub-task for $\textup{LLM}_{3}$.
Acc $\uparrow$ TPR $\uparrow$ FPR $\downarrow$
@@
MS 0.60 1.0 1.0
Sample-wise detection     We also use the different baselines for sample-wise detection of erroneous steps, as outlined in Section 3.3. A step is classified as incorrect if a baseline’s metric falls below a threshold. Table 1 presents the results for $\textup{LLM}_{3}$, with optimal thresholds chosen from a held-out dataset. Our method achieves significantly higher accuracy and fewer false positives than the baselines, making it more reliable for sample-wise error detection.
@@ 5.2 Arithmetic operations on Llama-3-8B
Following our toy experiments, we now evaluate our framework in a more realistic setting using the Llama-3-8B model [7]. We focus on a simple arithmetic task that involves both multiplication and addition. The goal is to assess the model’s performance on each operation.
Experimental setup     We sample two integers $x$ and $y$ uniformly from $[1,10^{5})$. The prompt to the model is:
@@
Model Accuracy     The model’s accuracy across steps is:
Step 1: 80%, Step 2: 98%, Step 3: 42%.
Most failures occur in Step 3, which involves adding previously computed values. Analyzing the $(x,y)$ distribution where the model is incorrect (Figure 4), we find that errors mainly arise when one variable is large and the other is small. This suggests that correctness depends heavily on $(x,y)$, making it difficult for baselines to pinpoint the erroneous step in the model’s CoT reasoning.
Figure 4: The distribution of $(x,y)$ for incorrect samples: Llama-3-8B struggles to add large and small numbers (represented by the top-left and bottom-right shaded regions).